# Random Feature Representation Boosting

### v1

@@ D.1 Preliminaries
=$\mathcal{H}=\{x\mapsto\Delta\tanh(Bx):\lambda_{\textnormal{max}}(\Delta)\textnormal{ and }\lambda_{\textnormal{max}}(B)\textnormal{ bounded by }\lambda\}.$
Here $\lambda_{\textnormal{max}}$ denotes the maximum singular value. We will use the properties that $\|\Delta\|_{2}\leq\lambda$, $\|\Delta_{k}\|_{1}\leq\lambda\sqrt{p}$, $\|\Delta_{k}\|_{\infty}\leq\lambda$, $\|B_{j}\|_{1}\leq\lambda\sqrt{D}$, and $\|B_{j}\|_{\infty}\leq\lambda$. Here $\Delta_{k}$ denotes the $k$-th row of $\Delta$, and $\|\cdot\|_{p}$ the $\ell_{p}$ vector norm or matrix spectral norm.
To prove Theorem 3.4, we need to compute the Rademacher complexities of RFRBoost, and verify that our class of weak learners satisfy all the assumptions of Theorem D.2. The only critical assumption to check is that $\sup_{g_{t}\in\mathcal{G}_{t},x\in\mathcal{X}}\|g_{t}(x)\|_{2}$ is bounded. This is the case for our particular model class, since $\|g_{t}(x)\|_{2}=\|\Delta_{t}\tanh{B_{t}\Phi_{t-1}(x)}\|_{2}\leq\|\Delta_{t}\|_{2}\|\tanh{B_{t}\Phi_{t-1}(x)}\|_{2}\leq\lambda\sqrt{p}$, which follows from basic properties of vector and matrix norms.
@@ D.2 RFRBoost Rademacher Computations
$=\sum_{k=1}^{D}\mathbb{E}_{\rho}\bigg{[}\sup_{\begin{subarray}{c}B\in\mathbb{R}^{p\times D}\\ \Delta\in\mathbb{R}^{D\times p}\\ \lambda_{\textnormal{max}}(\Delta),\lambda_{\textnormal{max}}(B)\leq\lambda_{1}\end{subarray}}\frac{1}{\widetilde{n}}\sum_{i=1}^{\widetilde{n}}\rho_{i,k}\langle\Delta_{k},\tanh(B\Phi_{t-1}(x_{t,i})\rangle\bigg{]}$
@@
Here we additionally used the fact that $\sup_{x}|g_{k}(x)|=\sup_{x}|\langle\Delta_{k},\tanh(B\Phi_{t-1})(x)\rangle|\leq\lambda_{1}\sqrt{p}$ for Lemma D.3. ∎
@@ E.2 Model Architectures and Hyperparameter Ranges for OpenML Tasks
We used the Aeon library (Middlehurst et al., 2024) to generate the critical difference diagrams of Section 4.1.
@@ E.3 Point Cloud Separation Task
For the point cloud separation task, we use the same model structure as in the OpenML experiments, but fix the number of ResNet layers to 3 residual blocks. We additionally fix the activation to $\tanh$ and feature dimension to 512 for all models. Further architectural details and learning rate annealing strategies are consistent with those described in Section 4.1 and Section E.2. The SWIM scale was fixed at 1.0 for all random feature models.
@@
Table 5: Hyperparameter Ranges for the point cloud separation task
@@
Logistic Regression l2 [1, 1e-1, 1e-2, 1e-3, 1e-4, 1e-5]
E2E MLP ResNet lr [1e-2, 1e-3, 1e-4, 1e-5, 1e-6]
@@
weight_decay $10^{-5}$
batch_size 64

### v2

@@ D.1 Preliminaries
=$\mathcal{H}=\{x\mapsto A\tanh(Bx):\lambda_{\textnormal{max}}(A)\textnormal{ and }\lambda_{\textnormal{max}}(B)\textnormal{ bounded by }\lambda\}.$
Here $\lambda_{\textnormal{max}}$ denotes the maximum singular value. We will use the properties that $\|A\|_{2}\leq\lambda$, $\|A_{k}\|_{1}\leq\lambda\sqrt{p}$, $\|A_{k}\|_{\infty}\leq\lambda$, $\|B_{j}\|_{1}\leq\lambda\sqrt{D}$, and $\|B_{j}\|_{\infty}\leq\lambda$. Here $A_{k}$ denotes the $k$-th row of $A$, and $\|\cdot\|_{p}$ the $\ell_{p}$ vector norm or matrix spectral norm.
To prove Theorem 3.4, we need to compute the Rademacher complexities of RFRBoost, and verify that our class of weak learners satisfy all the assumptions of Theorem D.2. The only critical assumption to check is that $\sup_{g_{t}\in\mathcal{G}_{t},x\in\mathcal{X}}\|g_{t}(x)\|_{2}$ is bounded. This is the case for our particular model class, since $\|g_{t}(x)\|_{2}=\|A_{t}\tanh{B_{t}\Phi_{t-1}(x)}\|_{2}\leq\|A_{t}\|_{2}\|\tanh{B_{t}\Phi_{t-1}(x)}\|_{2}\leq\lambda\sqrt{p}$, which follows from basic properties of vector and matrix norms.
@@ D.2 RFRBoost Rademacher Computations
$=\sum_{k=1}^{D}\mathbb{E}_{\rho}\bigg{[}\sup_{\begin{subarray}{c}B\in\mathbb{R}^{p\times D}\\ A\in\mathbb{R}^{D\times p}\\ \lambda_{\textnormal{max}}(A),\lambda_{\textnormal{max}}(B)\leq\lambda_{1}\end{subarray}}\frac{1}{\widetilde{n}}\sum_{i=1}^{\widetilde{n}}\rho_{i,k}\langle A_{k},\tanh(B\Phi_{t-1}(x_{t,i})\rangle\bigg{]}$
@@
Here we additionally used the fact that $\sup_{x}|g_{k}(x)|=\sup_{x}|\langle A_{k},\tanh(B\Phi_{t-1})(x)\rangle|\leq\lambda_{1}\sqrt{p}$ for Lemma D.3. ∎
@@ E.2 Model Architectures and Hyperparameter Ranges for OpenML Tasks
We used the Aeon library (Middlehurst et al., 2024) to generate the critical difference diagrams in Section 4.1. Below, we provide full dataset-wise results in Tables 5 to 7 for each model and dataset used in the evaluation, averaged across all 5 folds. Figures 5 and 6 visualize these results using scatter plots overlaid with box-and-whiskers plots, providing insight into the distribution and variability of performance across datasets. We refer the reader to the critical difference diagrams for a rigorous statistical comparison of model rankings across all datasets.
Table 5: Test classification scores on OpenML datasets.
Dataset ID Logistic Regression RFRBoost RFNN E2E MLP ResNet XGBoost
3 0.974 0.996 0.989 0.996 0.997
6 0.768 0.912 0.890 0.944 0.903
11 0.878 0.981 0.979 0.934 0.931
14 0.812 0.844 0.804 0.828 0.830
15 0.964 0.966 0.963 0.964 0.960
16 0.948 0.974 0.945 0.973 0.954
18 0.728 0.739 0.727 0.739 0.721
22 0.815 0.837 0.812 0.840 0.795
23 0.512 0.551 0.553 0.532 0.547
28 0.970 0.991 0.980 0.989 0.978
29 0.854 0.854 0.845 0.852 0.862
31 0.761 0.762 0.756 0.743 0.744
32 0.946 0.993 0.990 0.992 0.986
37 0.762 0.758 0.760 0.767 0.762
38 0.968 0.979 0.977 0.977 0.988
44 0.931 0.950 0.936 0.941 0.952
50 0.983 0.983 0.980 0.976 0.995
54 0.801 0.849 0.829 0.843 0.772
151 0.760 0.796 0.784 0.792 0.847
182 0.861 0.909 0.908 0.911 0.913
188 0.641 0.664 0.643 0.635 0.652
307 0.838 0.993 0.972 0.989 0.924
458 0.994 0.994 0.993 0.988 0.990
469 0.187 0.188 0.196 0.193 0.178
1049 0.908 0.910 0.904 0.907 0.919
1050 0.894 0.889 0.898 0.892 0.890
1053 0.808 0.807 0.806 0.810 0.809
1063 0.841 0.845 0.833 0.841 0.839
Table 6: Test classification scores on OpenML datasets.
Dataset ID Logistic Regression RFRBoost RFNN E2E MLP ResNet XGBoost
1067 0.853 0.855 0.858 0.855 0.853
1068 0.925 0.927 0.928 0.928 0.935
1461 0.902 0.902 0.904 0.900 0.903
1462 0.988 1.000 1.000 1.000 0.997
1464 0.766 0.791 0.794 0.769 0.783
1475 0.478 0.545 0.534 0.570 0.601
1480 0.715 0.703 0.703 0.671 0.684
1486 0.943 0.949 0.946 0.950 0.959
1487 0.935 0.945 0.945 0.944 0.940
1489 0.748 0.876 0.864 0.906 0.903
1494 0.873 0.882 0.868 0.878 0.873
1497 0.701 0.929 0.907 0.944 0.998
1510 0.977 0.974 0.974 0.961 0.961
1590 0.850 0.858 0.856 0.851 0.865
4534 0.938 0.952 0.949 0.960 0.956
4538 0.478 0.554 0.513 0.606 0.660
6332 0.735 0.791 0.763 0.791 0.796
23381 0.626 0.640 0.624 0.566 0.612
23517 0.505 0.516 0.502 0.489 0.500
40499 0.997 0.997 0.992 0.999 0.985
40668 0.752 0.775 0.759 0.777 0.800
40701 0.867 0.936 0.923 0.951 0.953
40966 0.994 0.999 0.989 0.997 0.984
40975 0.932 0.997 0.977 0.999 0.994
40982 0.718 0.752 0.758 0.755 0.800
40983 0.971 0.986 0.987 0.988 0.983
40984 0.901 0.934 0.930 0.927 0.929
40994 0.961 0.957 0.957 0.946 0.944
41027 0.685 0.812 0.798 0.844 0.838
Table 7: Test RMSE scores on OpenML regression datasets.
Dataset ID Ridge Regression Greedy RFRBoost $A_{scalar}$ Greedy RFRBoost $A_{diag}$ Greedy RFRBoost $A_{dense}$ Gradient RFRBoost RFNN E2E MLP ResNet XGBoost
41021 0.233 0.233 0.232 0.231 0.231 0.232 0.265 0.249
44956 0.647 0.631 0.621 0.615 0.616 0.622 0.614 0.633
44957 0.688 0.362 0.246 0.238 0.233 0.347 0.316 0.197
44958 0.595 0.239 0.205 0.193 0.191 0.246 0.246 0.039
44959 0.589 0.304 0.275 0.277 0.286 0.305 0.311 0.245
44960 0.290 0.069 0.050 0.049 0.049 0.088 0.200 0.037
44962 0.446 0.448 0.448 0.446 0.447 0.447 0.474 0.452
44963 0.841 0.784 0.773 0.749 0.752 0.783 0.703 0.694
44964 0.521 0.443 0.406 0.395 0.398 0.443 0.369 0.340
44965 0.883 0.867 0.873 0.860 0.858 0.871 0.864 0.839
44966 0.719 0.724 0.718 0.719 0.721 0.726 0.731 0.720
44967 0.786 0.788 0.790 0.788 0.784 0.795 0.814 0.789
44969 0.409 0.084 0.027 0.022 0.025 0.061 0.058 0.086
44970 0.645 0.609 0.610 0.618 0.615 0.603 0.616 0.622
44971 0.840 0.789 0.779 0.780 0.777 0.785 0.729 0.685
44972 0.796 0.776 0.775 0.768 0.767 0.774 0.777 0.722
44973 0.602 0.287 0.235 0.194 0.194 0.284 0.177 0.244
44974 0.457 0.214 0.184 0.153 0.152 0.211 0.158 0.126
44975 0.026 0.027 0.026 0.026 0.026 0.026 0.056 0.109
44976 0.270 0.192 0.177 0.170 0.170 0.196 0.161 0.192
44977 0.590 0.525 0.517 0.508 0.506 0.525 0.486 0.446
44978 0.242 0.149 0.145 0.146 0.148 0.149 0.142 0.134
44979 0.252 0.163 0.148 0.142 0.140 0.165 0.145 0.142
44980 0.766 0.430 0.324 0.318 0.317 0.438 0.275 0.457
44981 0.914 0.913 0.913 0.912 0.913 0.914 0.636 0.611
44983 0.431 0.263 0.250 0.247 0.244 0.257 0.235 0.229
44984 0.689 0.660 0.658 0.656 0.657 0.659 0.656 0.656
44987 0.334 0.233 0.194 0.183 0.192 0.246 0.229 0.211
44989 0.308 0.293 0.287 0.276 0.274 0.295 0.304 0.263
44990 0.404 0.395 0.396 0.395 0.393 0.399 0.405 0.399
44993 0.799 0.776 0.773 0.770 0.769 0.775 0.787 0.773
44994 0.265 0.215 0.216 0.215 0.216 0.216 0.231 0.225
45012 0.448 0.343 0.339 0.330 0.330 0.341 0.351 0.331
45402 0.621 0.526 0.507 0.499 0.497 0.532 0.474 0.502
Figure 5: Scatter and box-and-whiskers plot of classification accuracy across all OpenML datasets. Each dot corresponds to a single dataset, and box plots summarize the distribution for each model.
Figure 6: Scatter and box-and-whiskers plot of regression RMSE across all OpenML datasets. Each dot corresponds to a single dataset, and box plots summarize the distribution for each model.
@@ E.3 Point Cloud Separation Task
For the point cloud separation task, we use the same model structure as in the OpenML experiments, but fix the number of ResNet layers to 3 residual blocks. We additionally fix the activation to $\tanh$ and feature dimension to 512 for all models. Further architectural details and learning rate annealing strategies are consistent with those described in Section 4.1 and Section E.2. The SWIM scale was fixed at 1.0 for all random feature models. We use 5,000 randomly sampled data points for training and validation, and leave the remaining 5,000 for the final test set.
@@
Table 8: Hyperparameter ranges for the point cloud separation task.
@@
Logistic Regression l2 [1, 1e-1, 1e-2, 1e-3, 1e-4]
E2E MLP ResNet lr [1e-1, 1e-2, 1e-3, 1e-4, 1e-5]
@@
weight_decay 1e-5
batch_size 128
E.4 SWIM vs i.i.d. Ablation Study
In this section, we present a small ablation study comparing SWIM random features to standard i.i.d. Gaussian random features across OpenML regression and classification tasks. The evaluation procedure is the same as presented in Section 4.1, the only change being that the SWIM scale parameter is replaced by an i.i.d. scale parameter, ranging logarithmically from 0.1 to 10 for all models. Pairwise comparisons for each model variant are visualized in Figures 7 to 8, where each point represents a dataset and compares the performance of SWIM and i.i.d. features. Summary statistics, including average RMSE/accuracy, number of dataset-wise wins, and average training time, are reported in Tables 9 to 10.
Overall, performance differences between SWIM and i.i.d. features are small but consistent. For regression, SWIM-based models achieve lower mean RMSE across all variants, and win on 22 out of 34 datasets in the best case (Greedy RFRBoost $A_{\text{dense}}$). For classification, accuracy is nearly identical across methods, though i.i.d. features slightly outperform SWIM in the total number of wins (32 out of 56). SWIM-based methods typically incur slightly higher training time, reflecting the additional structure in their initialization. These results suggest that while SWIM can offer marginal gains in accuracy or RMSE, the benefits must be balanced against training efficiency.
Table 9: Comparison of SWIM vs i.i.d. feature initialization on OpenML regression tasks.
Model Variant Mean RMSE Wins Fit Time (s)
RFNN SWIM 0.434 20 0.053
i.i.d. 0.441 14 0.029
Gradient RFRBoost SWIM 0.408 18 1.688
i.i.d. 0.415 16 1.816
Greedy RFRBoost $A_{dense}$ SWIM 0.408 22 2.734
i.i.d. 0.416 12 2.667
Greedy RFRBoost $A_{diag}$ SWIM 0.415 19 1.631
i.i.d. 0.415 15 1.490
Greedy RFRBoost $A_{scalar}$ SWIM 0.434 22 1.024
i.i.d. 0.443 12 0.720
Table 10: Comparison of SWIM vs i.i.d. feature initialization on OpenML classification tasks.
Model Variant Mean Accuracy Wins Fit Time (s)
RFNN SWIM 0.845 25 1.189
i.i.d. 0.844 31 0.634
Gradient RFRBoost SWIM 0.853 32 2.519
i.i.d. 0.850 24 2.350
Figure 7: Comparison of classification accuracy using SWIM versus i.i.d. Gaussian random features across OpenML datasets. Each scatter plot compares accuracy on a per-dataset basis for different model variants. Points below the diagonal indicate better performance with SWIM.
Figure 8: Comparison of RMSE using SWIM versus i.i.d. Gaussian random features across OpenML regression tasks. Each scatter plot compares RMSE per dataset for different model variants. Points above the diagonal indicate better performance with SWIM.
E.5 Supplementary Larger-Scale Experiments
To further assess the scalability and performance of RFRBoost relative to baseline models, we conducted experiments on four larger datasets: two of the largest from the OpenML Benchmark suite used in this work (each with approximately 100k samples — one classification: ID 23517; one regression: ID 44975), and two widely-used benchmark datasets with approximately 500k samples (one classification: CoverType (Blackard, 1998); one regression: YearPredictionMSD (YPMSD) (Bertin-Mahieux, 2011)).
For these larger-scale evaluations, we employed a hold-out test set strategy. Hyperparameters for each model were selected via a grid search performed on a 20% validation split of the training data. The models considered are consistent with those in Section 4.1: E2E MLP ResNet, Gradient RFRBoost, Logistic/Ridge Regression, RFNN, and XGBoost. To analyze performance trends with increasing data availability, models were trained on exponentially increasing subsets of the training data, starting from 2048 samples ($2^{11}$) up to the maximum available training size for each dataset.
For the OpenML datasets, the test sets were defined as 33% (ID 23517, classification) and 11% (ID 44975, regression) of the total instances, resulting in a maximum training set size of approximately 64,000 instances ($2^{16}$). For YPMSD (regression), we adhered to the designated split: the first 463,715 examples for training and the subsequent 51,630 examples for testing. For CoverType (classification), we used the standard train-test split provided by the popular Python machine learning library Scikit-learn (Pedregosa et al., 2011), yielding 464,809 training and 116,203 test instances.
The hyperparameter grids used for each model in these experiments are detailed in Table 11. To account for variability, each experiment was repeated 5 times with different random seeds for each model. All experiments were carried out on a single NVIDIA RTX 6000 (Turing architecture) GPU.
Table 11: Hyperparameter grid for the larger-scale experiments.
Model Hyperparameter Values
RFRBoost l2_cls [1e-1, 1e-2, 1e-3, 1e-4, 1e-5]
n_layers [1, 3, 6]
feature_dim 512
boost_lr 1.0
RFNN l2 [1e-1, 1e-2, 1e-3, 1e-4, 1e-5]
feature_dim 512
Logistic/Ridge Regression l2 [1e-1, 1e-2, 1e-3, 1e-4]
XGBoost lr [0.1, 0.033, 0.01]
n_estimators [250, 500, 1000]
max_depth [1, 3, 6, 10]
lambda 1.0
E2E MLP ResNet lr [1e-1, 1e-2, 1e-3]
n_epochs [10, 20, 30]
feature_dim 512
end_lr_factor 0.01
weight_decay 1e-5
batch_size 256
The model training times and predictive performance (RMSE for regression, accuracy for classification) as a function of training set size are presented in Figure 9 and Figure 10, respectively. Regarding training times shown in Figure 9, Ridge/Logistic Regression and RFNNs consistently exhibit the fastest training, as expected. RFRBoost consistently trains faster than both XGBoost and the E2E MLP ResNets across all datasets. Notably, on the OpenML ID 23517 dataset, RFRBoost trains faster than the single-layer RFNN. This may be attributed to RFNNs fitting a potentially very wide random feature layer directly to the output targets (logits or regression values), which can be computationally intensive. In contrast, RFRBoost constructs a comparatively shallow representation layer-wise with random feature residual blocks, where only the final, potentially lower-dimensional, representation is mapped to the output targets. In some scenarios, this may lead to lower computational times.
Considering the predictive performance shown in Figure 10, RFRBoost consistently and significantly outperforms the baseline single-layer RFNN across all datasets and training sizes, demonstrating its effectiveness in using depth for random feature models to increase performance. This improvement is particularly pronounced on the larger CoverType and YPMSD datasets, underscoring the benefit of RFRBoost’s deep, layer-wise construction. However, in the larger data regimes E2E trained MLP ResNets and XGBoost generally achieve superior predictive performance. The OpenML ID 23517 dataset stands out, however, as RFRBoost and Logistic Regression not only exhibit lower variance but also achieve the best overall performance. Furthermore, we find that XGBoost significantly underperforms on OpenML ID 44975 compared to all other baseline models. Given RFRBoost’s strong performance relative to its computational cost and its ability to construct meaningful deep representations, exploring its use as an initialization strategy for subsequent fine-tuning of end-to-end trained networks presents a promising avenue for future research, potentially further enhancing the performance of E2E models.
Figure 9: Model training time (seconds, log scale) versus training set size on larger datasets. Lines represent the mean training time over 5 runs, and shaded areas indicate 95% confidence intervals.
Figure 10: Model predictive performance (RMSE for regression, accuracy for classification) versus training set size on larger datasets. Lines represent the mean performance over 5 runs, and shaded areas indicate 95% confidence intervals.